OpenStudy (anonymous):

This formula converts Fahrenheit to Celsius. C = 5 - 9 (F−32)

OpenStudy (kainui):

Ok thanks.

OpenStudy (anonymous):

for what

OpenStudy (anonymous):

saying that \(C=\frac{5}{9}(F-32)\)

OpenStudy (anonymous):

but why would he say thanks

OpenStudy (anonymous):

he was being a wise guy you wrote a formula with no question

OpenStudy (anonymous):

oh whoops!Which is a solution for F?

OpenStudy (anonymous):

i need helphelphel

OpenStudy (shane_b):

Here are the steps: \[C=\frac{5}{9}(F-32)\]\[\frac{9}{5}C=F-32\]\[\frac{9}{5}C+32=F\]

OpenStudy (anonymous):

what do i do now

OpenStudy (anonymous):

someone please help

OpenStudy (shane_b):

You asked for the solution for F which is the last line of my last post.